### v1

@@ 3 Active Learning in Bird Sound Classification
Challenges and Motivation. In PAM, a feature vector $\mathbf{x}\in\mathcal{X}$ represents a $D$-dimensional instance, originating from either a focal recording where $\mathcal{X}=\mathcal{F}$, or a soundscape recording with $\mathcal{X}=\mathcal{S}$. Focal recordings are extensively available on the citizen-science platform Xeno-Canto (XC) [10] with a global collection of over 800,000 recordings, making them particularly suitable for model training. Large-scale bird sound classification models (e.g., BirdNET[2]) are primarily trained on focals. These multi-class recordings feature isolated bird vocalizations where each instance $\mathbf{x}$ is associated with a class label $y\in\mathcal{Y}$, where $\mathcal{Y}=\{1,...,C\}$. The focal data distribution is denoted as $p_{\texttt{Focal}}(\mathbf{x},y)$. However, annotations from XC often come with weak labels, lacking precise vocalization timestamps. As noted by Van Merriënboer et al. [11]. As noted by Van Merriënboer et al. [11], evaluating on focals does not adequately reflect a model’s generalization performance in real-world PAM scenarios, rendering them unsuitable for assessing deployment capabilities. Soundscape recordings are passively recorded in specific regions, capturing the entire acoustic environment for PAM projects using static ARUs over extended periods. For instance, the High Sierra Nevada (HSN) [2] dataset includes long-duration soundscapes with precise labels and timestamps from multiple sites. These recordings are treated as multi-label tasks and are valuable for assessing model deployment in real-world PAM. Each instance $\mathbf{x}$ is associated with multiple class labels $y\in\mathcal{Y}$, represented by a one-hot encoded multi-label vector $\mathbf{y}=[y_{1},\ldots,y_{C}]\in{[0,1]}^{C}$. An instance can contain no bird sounds, represented by a zero-vector $\mathbf{y}=\mathbf{0}\in\mathbb{R}^{C}$. Soundscapes’ limited scale and the extensive annotation effort make them less suitable for large-scale model training. Yet, we believe that they are ideal for fine-tuning and adaptation in practical environments. We denote the soundscape data distribution as $p_{\texttt{Scape}}(\mathbf{x},\mathbf{y})$. The disparity in data distributions, $p_{\texttt{Scape}}(\mathbf{x},\mathbf{y})\neq p_{\texttt{Focal}}(\mathbf{x},y)$, leads to a distribution shift that impacts the performance of SOTA bioacoustic models trained on focals when deployed in PAM, where only a few labeled soundscapes are available for training. Therefore, we propose using deep AL to efficiently adapt models to PAM scenarios.
@@ 4 Experiments
Setup. We employ Google’s Perch as the pre-trained feature extractor with a feature dimensionality of $D=1280$, following Ghani et al. [3]. Each iteration of the AL cycle involves initializing and training the last DNN layer for 200 epochs using the Rectified Adam optimizer [12] (batch size: 128, learning rate: 0.05, weight decay: 0.0001) with a cosine annealing scheduler [13]. The hyperparameters are empirically determined with convergence on random train samples as done in [14]. We utilize the HSN dataset [15] from BirdSet [6], consisting of $5,280$ 5-second soundscape segments from the initial three days of recordings for our unlabeled pool, and $6,720$ segments from the last two days for testing. Initially, $10$ instances are selected randomly, followed by 50 iterations of $b{=}10$ acquisitions each, totaling a budget of ${B}{=}510$. We benchmark against Random acquisitions and use Typiclust [16] and Badge[17] as diversity-based and hybrid strategies, respectively. As an uncertainty-based strategy, we employ the mean Entropy of all binary predictions. The effectiveness of each strategy is assessed by analyzing the learning curves through a collection of threshold-free metrics [6]: T1-Acc., class-based mean average precision (cmAP), and area under the receiver operating characteristic curve (AUROC). The metrics are computed on the test dataset post-training in each cycle, with learning curve improvements averaged over ten repetitions for consistency.
@@
Results. We present the improvement curves for the metric collection in Figure 2. The results demonstrate that no single strategy is universally superior across all metrics. However, nearly all metrics show enhanced performance compared to Random. Notably, Typiclust displays strong performance across all metrics at the start of the deep AL cycle, supporting the findings of [16] that a diverse selection is beneficial at the cycle’s onset. However, its effectiveness diminishes over time when diversity becomes less crucial. Conversely, except for the AUROC metric where Entropy initially performs poorly but strongly improves over time, Entropy outperforms in all iterations for cmAP and T1-Acc, showing a consistent improvement over Random improvement of up to 15%.
@@
Figure 2: Improvement curves of deep AL selection strategies Badge, Entropy, and Typiclust over Random with the metric collection a) AUROC, b) cmAP and c) T1-Acc. The results are averaged over ten randomly initialized repetitions to ensure consistency.
@@ 5 Conclusion
In this work, we demonstrated the potential of deep active learning (AL) in computational avian bioacoustics. We showed how deep AL can be integrated into real-world passive acoustic monitoring by utilizing BirdSet, where a rapid model adaption through fine-tuning on soundscape recordings is advantageous for the identification of bird species. Our results indicate that employing selection strategies in deep AL enhances model performance and accelerates adaptation compared to random sampling. For future work, we aim to expand the implementation of deep AL in avian bioacoustics utilizing all datasets from the BirdSet dataset collection to provide more robust performance insights and explore more advanced query strategies [13, 18].

### v2

@@ 3 Active Learning in Bird Sound Classification
Motivation. In PAM, a feature vector $\mathbf{x}\in\mathcal{X}$ represents a $D$-dimensional instance, originating from either a focal recording where $\mathcal{X}=\mathcal{F}$, or a soundscape recording with $\mathcal{X}=\mathcal{S}$. Focal recordings are extensively available on the citizen-science platform Xeno-Canto (XC) [10] with a global collection of over 800,000 recordings, making them particularly suitable for model training. Large-scale bird sound classification models (e.g., BirdNET[2]) are primarily trained on focals. These multi-class recordings feature isolated bird vocalizations where each instance $\mathbf{x}$ is associated with a class label $y\in\mathcal{Y}$, where $\mathcal{Y}=\{1,...,C\}$. The focal data distribution is denoted as $p_{\texttt{Focal}}(\mathbf{x},y)$. However, annotations from XC often come with weak labels, lacking precise vocalization timestamps. As noted by Van Merriënboer et al. [11], evaluating on focals does not adequately reflect a model’s generalization performance in real-world PAM scenarios, rendering them unsuitable for assessing deployment capabilities. Soundscape recordings are passively recorded in specific regions, capturing the entire acoustic environment for PAM projects using static ARUs over extended periods. For instance, the High Sierra Nevada (HSN) [2] dataset includes long-duration soundscapes with precise labels and timestamps from multiple recording sites. Soundscapes are treated as multi-label tasks and are valuable for assessing model deployment in real-world PAM. Each instance $\mathbf{x}$ is associated with multiple class labels $y\in\mathcal{Y}$, represented by a one-hot encoded multi-label vector $\mathbf{y}=[y_{1},\ldots,y_{C}]\in{[0,1]}^{C}$. An instance can contain no bird sounds, represented by a zero-vector $\mathbf{y}=\mathbf{0}\in\mathbb{R}^{C}$. Soundscapes’ limited scale and the extensive annotation effort make them less suitable for large-scale model training. We denote the soundscape data distribution as $p_{\texttt{Scape}}(\mathbf{x},\mathbf{y})$. The disparity in data distributions, $p_{\texttt{Scape}}(\mathbf{x},\mathbf{y})\neq p_{\texttt{Focal}}(\mathbf{x},y)$, leads to a distribution shift that impacts the performance of SOTA bioacoustic models trained on focals when deployed in PAM. Additionally, highly diverse deployment conditions in PAM projects - such as background noise, recording devices, and their locations - also lead to domain differences within and between soundscape recordings. These variations further highlight the need for compact models that can quickly and easily adapt to changing environments. Thus, we argue that using labeled soundscapes in novel deployment scenarios for fine-tuning the model is vital. Therefore, we propose deep AL to enable fast model adaption to various PAM scenarios.
@@ 4 Experiments
Setup. We employ Google’s Perch as the pre-trained feature extractor with a feature dimensionality of $D=1280$, following Ghani et al. [3]. Each iteration of the AL cycle involves initializing and training the last DNN layer for 200 epochs using the Rectified Adam optimizer [12] (batch size: 128, learning rate: 0.05, weight decay: 0.0001) with a cosine annealing scheduler [13]. The hyperparameters are empirically determined with convergence on random train samples as done in [14]. We utilize the HSN dataset [15] from BirdSet [6], consisting of $5,280$ 5-second soundscape segments from the initial three days of recordings for our unlabeled pool. Thus, we simulate practical deployment scenario where we initially collect data from various recording sites that we want to quickly adapt the model to and reduce annotation effort. Subsequently, we utilize $6,720$ segments from the last two days for testing model performance. Initially, $10$ instances are selected randomly, followed by 50 iterations of $b{=}10$ acquisitions each, totaling a budget of ${B}{=}510$. We benchmark against Random acquisitions and use Typiclust [16] and Badge[17] as diversity-based and hybrid strategies, respectively. As an uncertainty-based strategy, we employ the mean Entropy of all binary predictions. The effectiveness of each strategy is assessed by analyzing the learning curves through a collection of threshold-free metrics [6]: T1-accuracy, class-based mean average precision (cmAP), and area under the receiver operating characteristic curve (AUROC). The metrics are computed on the test dataset post-training in each cycle, with learning curve improvements averaged over ten repetitions for consistency.
@@
Results. We present the improvement curves for the metric collection in Figure 2. The results demonstrate that no single strategy is universally superior across all metrics. However, nearly all metrics show enhanced performance compared to Random. Notably, Typiclust displays strong performance across all metrics at the start of the deep AL cycle, supporting the findings of [16] that a diverse selection is beneficial at the cycle’s onset. However, its effectiveness diminishes over time when diversity becomes less crucial. Conversely, except for the AUROC metric where Entropy initially performs poorly but strongly improves over time, Entropy outperforms in all iterations for cmAP and T1-Acc, showing a consistent improvement over Random of up to 15%.
@@
Figure 2: Improvement curves of deep AL selection strategies Badge, Entropy, and Typiclust over Random with the metric collection a) AUROC, b) cmAP and c) T1-Acc. The results are averaged over ten randomly initialized repetitions to ensure consistency and the standard deviation is displayed.
5 Open Challenges and Limitations
This pilot study explores the use of deep AL to tailor avian bioacoustic models for various deployment scenarios in PAM. Although the initial results are encouraging, they remain preliminary. Several key challenges, which are outlined below, need to be addressed to fully realize the potential of deep AL in this field.
Pool creation. The limited availability of soundscape data, which is primarily used for model evaluation [6], poses challenges in creating pool datasets for deep AL. The process of generating a fine-tuning training pool can affect class balance and raises concerns about the composition methodology. Additionally, in scenarios where data are sourced from PAM projects, the variability in recording sites is often not disclosed in publicly available datasets. This lack of information makes it challenging to create a diverse and representative training pool that takes recording locations into account. To effectively investigate deep AL, a transparent approach to dataset generation is essential.
Deployment in practice. Deploying deep AL in real-world PAM environments requires addressing several practical considerations. These include determining optimal batch sizes for data annotation and effectively allocating the total budget. The labor-intensive and costly process of labeling PAM recordings, which requires human expertise [18], highlights the need for accurately estimating the expected annotation effort. Additionally, exploring various deployment settings and tasks can reveal the versatility and potential challenges of applying deep AL, leading to more effective and scalable solutions for avian bioacoustics. For instance, tasks might involve not only classifying bird species but also identifying specific call densities [19], which would require modifications to the model evaluation process.
Evaluation. Traditional metrics such as AUROC, cmAP, and T1-Acc offer a general overview of model performance but may be inadequate in practice-specific scenarios, such as ensuring a high recall of a specific species or identifying bird call density [19]. A more nuanced approach to evaluating deep AL models involves customizing metrics to align with practical objectives, such as consistently identifying specific species. Enhancing evaluation methodologies to capture these specialized requirements is crucial for advancing the effectiveness of deep AL in real-world PAM applications.
@@ 6 Conclusion
In this work, we demonstrated the potential of deep active learning (AL) in computational avian bioacoustics. We showed how deep AL can be integrated into real-world passive acoustic monitoring by utilizing BirdSet, where a rapid model adaption through fine-tuning on soundscape recordings is advantageous for the identification of bird species. Our results indicate that employing selection strategies in deep AL enhances model performance and accelerates adaptation compared to random sampling. For future work, we aim to expand the implementation of deep AL in avian bioacoustics utilizing all datasets from the BirdSet dataset collection to provide more robust performance insights and explore additional query strategies [13, 20].